OpenStudy (anonymous):

(sin(theta/2)+cos(theta/2))^2 = 1+sin(x)

OpenStudy (anonymous):

establish identity

OpenStudy (anonymous):

expand the equation and use pythagoraus theorem

OpenStudy (campbell_st):

ok... so expand the left hand side \[\sin^2(\frac{\theta}{2}) + 2\sin(\frac{\theta}{2})\cos(\frac{\theta}{2}) + \cos^2(\frac{\theta}{2})\] now sin^2 + cos^2 = 1 so you problem is now \[1 + 2\sin(\frac{\theta}{2})\cos(\frac{\theta}{2})\] and you should know sin(2a) = 2sin(a)cos(a) apply this trig identity to the simplification above. hope it helps

OpenStudy (campbell_st):

think about a = @/2